# Fantastyc: Blockchain-based Federated Learning Made Secure and Practical

### v1

@@ I Introduction
  • Efficient Resource Use and Proofs: Our solution optimizes blockchain usage by storing updates and models within an external distributed storage connected to the blockchain. In this setup, the blockchain exclusively holds a cryptographic fingerprint of the data (the anchor), accompanied by a validity proof. This proof certifies the availability and integrity of the data within the storage. This approach ensures that clients can always retrieve computed models (availability) providing them with a guarantee of retrieving a trustworthy model (integrity) [^1] It is important to note that prevalent architectures in the blockchain space, utilizing an external storage system such as IPFS, often lack both computational integrity assurance and availability. These architectures solely offer tamper-proof guarantees: clients can only ascertain whether someone has tampered with their own data, without ensuring that data are the output of correct computations..
@@ IV-A2 Byzantine Servers
At high level, the protocol to generate the PoA&\&&I works as follows: as soon as a server gathers a value $v$ for which a PoA&\&&I is required, the server will verify if the value is valid. If valid, it stores the pair $(hash(v),v)$ in the key-value store where the hash of $v$ is used as key, and then broadcasts a signed message with the hash of the valid value, called the hash-key, to the other servers. As soon as a server receives a $f_{s}+1$ quorum of signed valid hash-keys from other servers, it produces a PoA&\&&I333 PoA&\&&I is similar in spirit to the Proof of Availability and Retrieval [21], which, however, does not target computational integrity. Indeed, [21] describes a mechanism to produce a proof of availability to implement a secure pull communication mechanism for blocks inside blockchains.. In the first step of the protocol, the verification performed to validate values depends on the type of value: in the case $v$ is a client update, the server verifies if the update comes from the set of clients selected for the current round; if $v$ is a model aggregated by other servers, then the value is considered valid if the received model and the locally computed model are not significantly different, typically within an epsilon distance.
@@ V System Detailed Design
In this section we first introduce the system model and a formalisation of the PoA&\&&I, before presenting the workflow of the FL process, which includes algorithms to build the PoA&\&&I (Section V-A) and methods to manage the non-determinism of the computations (Section V-A2). Finally, we defer the correctness arguments to Appendix B.
@@ V-A Solution workflow and algorithms
    Once the $fl_{i}$ task is registered on the blockchain, all clients and servers are aware of it. Each client sends a transaction to the blockchain to state its interest to join (if any). Such a transaction stakes a certain amount of coins666“Staking” in the context of cryptocurrency refers to the process of actively participating in the operations of a blockchain network. This involves locking up a certain amount of cryptocurrency as collateral to perform these tasks. for $fl_{i}$, witnessing the client commitment to participate. When the client’s transaction is registered on the blockchain we say that the client is registered for $fl_{i}$. Servers wait for sufficiently many clients to be registered (the number of awaited clients is set in $param_{fl_{i}}$).
@@
    When a server collects sufficiently many updates’s PoA&\&&I for round $t$ (as determined in $param_{fl_{i}}$), it has a candidate update set, $\mathcal{P}(G_{t})_{cand}$. It computes $C_{t+1}$, the clients subset to perform the learning in the next round and build $P(C_{t+1})$. Finally, it sends a transaction ${\tt NR}$ (New Round) with $\mathcal{P}(G_{t})_{cand}$, $C_{t+1}$ and $C_{t+1}$;
@@
Interestingly, as we detailed, for each round $t\geq 1$ the blockchain registers a single transaction per round. [^8] In the practice there is one transaction per server, but processes wait only for the first valid one on-chain to proceed with their computation.
@@ V-A2 Floating points and non-determinism handling
On a positive note, the disparity between computed models is small enough (cf. note 2) to introduce an acceptance vector gap that we denote as $\epsilon\in\Re^{d}$. We assume that $\epsilon$ is static and provided by the model owner when specifying $fl_{i}$. We assume that $\epsilon$ is sufficiently large to prevent a deadlock: a round in which there are no models issued by servers that obtain a PoA&\&&I (if their difference is bigger than $\epsilon$).
@@ VI Experimental Evaluation
It is worth noting that our evaluation does not deal with the demonstration of high throughput because we focus on only one FL task. Let us also stress that FL operates in rounds that can be lengthy, not imposing strict latency constraints.
@@ VI-E Scalability evaluation
TABLE II: Communication overhead with different fault thresholds
@@ VII Conclusion and Perspectives
Appendix A Solution Sequence Diagram
Fig. 5 details the workflow described in Section V, involving replicated servers, storage replicas, a blockchain, clients, and a model owner. In particular, we use a replicated activation box image on the server lifeline, corresponding to the build PoA&\&&I operation. With some abuse of notation, this represents that this operation involves all servers together.
Figure 5: Sequence diagram of the workflow described in Section V.
Appendix B Correctness Arguments
In the following we discuss the correctness argument related to the solution detailed in Section V.
Lemma 1
For any $({{\tt TAG}},v)$, if exists a valid $P({{\tt TAG}},v)$, then ${{\sf locally\_valid}}({{\tt TAG}},v)$ holds for at least one correct server, and eventually, $v$ is available from the distributed storage.
Proof 1 (Sketch)
In the next we consider the following Assumption:
Assumption 2
Lemma 2
Proof 2 (Sketch)
We start from round $t=0$:
Theorem 1
Proof 3 (Sketch)
Theorem 2
Proof 4 (Sketch)

### v2

@@ I Introduction
  • Efficient Resource Use and Proofs: Our solution optimizes blockchain usage by storing updates and models within an external distributed storage connected to the blockchain. In this setup, the blockchain exclusively holds a cryptographic fingerprint of the data (the anchor), accompanied by a validity proof. This proof certifies the availability and integrity of the data within the storage. This approach ensures that clients can always retrieve computed models (availability) providing them with a guarantee of retrieving a trustworthy model (integrity). It is important to note that prevalent architectures in the blockchain space, such as those utilizing external storage systems like IPFS, lack computational integrity even when they offer mechanisms for making redundant copies (proof-of-replication [13]). Indeed, the primary goal of these architectures is to offer tamper-proof guarantees: clients can only ascertain whether someone has tampered with their own data, without ensuring that the data are the output of correct computations.
@@ IV-A2 Byzantine Servers
At high level, the protocol to generate the PoA&\&&I works as follows: as soon as a server gathers a value $v$ for which a PoA&\&&I is required, the server will verify if the value is valid. If valid, it stores the pair $(hash(v),v)$ in the key-value store where the hash of $v$ is used as key, and then broadcasts a signed message with the hash of the valid value, called the hash-key, to the other servers. As soon as a server receives a $f_{s}+1$ quorum of signed valid hash-keys from other servers, it produces a PoA&\&&I222 PoA&\&&I is similar in spirit to the Proof of Availability and Retrieval [22], which, however, does not target computational integrity. Indeed, [22] describes a mechanism to produce a proof of availability to implement a secure pull communication mechanism for blocks inside blockchains.. In the first step of the protocol, the verification performed to validate values depends on the type of value: in the case $v$ is a client update, the server verifies if the update comes from the set of clients selected for the current round; if $v$ is a model aggregated by other servers, then the value is considered valid if the received model and the locally computed model are not significantly different, typically within an epsilon distance.
@@ V System Detailed Design
In this section we first introduce the system model and a formalisation of the PoA&\&&I, before presenting the workflow of the FL process, which includes algorithms to build the PoA&\&&I (Section V-A) and methods to manage the non-determinism of the computations (Section V-A2). We finally present correctness arguments of the solution (Section V-B).
@@ V-A Solution workflow and algorithms
    Once the $fl_{i}$ task is registered on the blockchain, all clients and servers are aware of it. Each client sends a transaction to the blockchain to state its interest to join (if any). Such a transaction stakes a certain amount of coins555“Staking” in the context of cryptocurrency refers to the process of actively participating in the operations of a blockchain network. This involves locking up a certain amount of cryptocurrency as collateral to perform these tasks. for $fl_{i}$, witnessing the client commitment to participate. When the client’s transaction is registered on the blockchain we say that the client is registered for $fl_{i}$. Servers wait for sufficiently many clients to be registered (the number of awaited clients is set in $param_{fl_{i}}$).
@@
    When a server collects sufficiently many updates’s PoA&\&&I for round $t$ (as determined in $param_{fl_{i}}$), it has a candidate update set, $\mathcal{P}(G_{t})_{cand}$. It computes $C_{t+1}$, the clients subset to perform the learning in the next round and build $P(C_{t+1})$. Finally, it sends a transaction ${\tt NR}$ (New Round) with $\mathcal{P}(G_{t})_{cand}$, $C_{t+1}$ and $P(C_{t+1})$;
@@
Interestingly, as we detailed, for each round $t\geq 1$ the blockchain registers a single transaction per round. [^7] In the practice there is one transaction per server, but processes wait only for the first valid one on-chain to proceed with their computation. Fig. 2 details the workflow, involving replicated servers, storage replicas, a blockchain, clients, and a model owner. In particular, we use a replicated activation box image on the server lifeline, corresponding to the build PoA&\&&I operation. With some abuse of notation, this represents that this operation involves all servers together.
Figure 2: Sequence diagram of the Fantastyc workflow.
@@ V-A2 Floating points and non-determinism handling
On a positive note, the disparity between computed models is small enough (cf. note 1) to introduce an acceptance vector gap that we denote as $\epsilon\in\Re^{d}$. We assume that $\epsilon$ is static and provided by the model owner when specifying $fl_{i}$. We assume that $\epsilon$ is sufficiently large to prevent a deadlock: a round in which there are no models issued by servers that obtain a PoA&\&&I (if their difference is bigger than $\epsilon$).
@@
V-B Correctness arguments
In the next we consider the following Assumption:
Assumption 2
With this assumption, we abstract the implementation of the selection function, assuming only that the selection is verifiable. In practice, the selection function can be secret to ensure resilience against an adaptive adversary that can choose which clients to corrupt, or it can be public if only a static adversary is assumed.
V-B1 Safety
Lemma 1
(availability and integrity) For any $({{\tt TAG}},v)$, if exists a valid $P({{\tt TAG}},v)$, then ${{\sf locally\_valid}}({{\tt TAG}},v)$ holds for at least one correct server, and eventually, $v$ is available from the distributed storage.
Proof 1 (Sketch)
Theorem 1
Proof 2 (Sketch)
V-B2 Liveness
Lemma 2
Proof 3 (Sketch)
We start from round $t=0$:
Theorem 2
Proof 4 (Sketch)
@@ VI Experimental Evaluation
It is worth noting that our evaluation does not deal with the demonstration of high throughput because we focus on only one FL task. Let us also stress that FL operates in rounds that can be lengthy, not imposing strict latency constraints. Furthermore, we do not compare our approach with other blockchain-based methods due significant differences in architectures, use and choice of blockchains, along with the impracticality of storing large models in the blockchain (see Section III for a discussion on related work)– an issue that limits realistic implementations.
@@ VI-E Scalability evaluation
TABLE II: Communication overhead with different fault thresholds.
@@ VII Conclusion and Perspectives
Acknowledgements
This work is funded by The Fantastyc Carnot project.